OpenStudy (anonymous):

HELP NEEDED, PLEASE! On one morning, communicating to work, Golda travels 10 miles in about 15 minutes and Dale travels 23 miles in about 30 miles. who has the fast average speed? Explain your answer.

OpenStudy (anonymous):

A similar was asked yesterday 10/4 this is the first one 23/2 for second one second > first Dale is fast

OpenStudy (anonymous):

I was right then. I was confused that's why asked. Thanks Anyways! :)

OpenStudy (anonymous):

No wait, from where did you get 10/(4) ? and 23/ (2) ??

OpenStudy (anonymous):

and there's a mistake in the question that is, HELP NEEDED, PLEASE! On one morning, communicating to work, Golda travels 10 miles in about 15 minutes and Dale travels 23 miles in about 30 minutes. who has the fast average speed? Explain your answer.

OpenStudy (anonymous):

You can calculate this in miles per minute or miles per hour. To calculate miles per minute you divide the miles by the minutes. To calculate the miles per hour you need to calculate how many fractions of an hour you have and multiply by the inverse. So 15 is 1/4 of an hour so you multipl the miles by 4 to get the miles per hour. 30 minutes is 1/2 an hour so you multiply by 2. So for miles per hour the answers are: \[10*4=40\frac{miles}{hour}\] \[23*2=46\frac{miles}{hour}\] For minutes we end up with decimals or fractions: \[\frac{10}{15}=\frac{2}{3}\frac{miles}{minute}\] \[\frac{23}{30}\frac{miles}{minute}\]